# Privacy Risks and Preservation Methods in Explainable Artificial Intelligence: A Scoping Review

### v2

@@ 4.2.4 Property inference
This type of risk pertains to inference of properties from the training data such as global statistics or aggregates (Mahloujifar et al., 2022), which model owners did not intend on sharing (Ganju et al., 2018) (Figure 9). The inferred property may not correspond to features in the training data or be correlated to any feature. A property inference model that uses the model outputs, $f(x)$, to infer such property, $p$, can be expressed as the following:
@@
Membership inference Interpretable Decision tree Naretto et al. (2022)
Interpretable (surrogate) Trepan Naretto et al. (2022)
Example-based Influence functions Shokri et al. (2020; 2021)
Counterfactuals Kuppa & Le-Khac (2021); Pawelczyk et al. (2023)
Self-influence functions Cohen & Giryes (2024)
Feature-based Gradient, integrated gradients, guided backpropagation, LRP, LIME, SmoothGrad Shokri et al. (2020; 2021)
Integrated gradients, SmoothGrad, VarGrad, Grad-CAM++, SHAP, LIME Liu et al. (2024)
Shapley values Ma et al. (2024)
Model inversion Interpretable Decision tree, rule list Ferry et al. (2024)
Example-based Influence functions Shokri et al. (2020; 2021)
Counterfactuals (native) (Goethals et al., 2023)
Feature-based Gradient, gradient x input, class activation maps (CAM), Grad-CAM, LRP (Zhao et al., 2021)
Integrated gradients, DeepLIFT, GradientSHAP, SmoothGrad Duddu & Boutet (2022)
Shapley values Luo et al. (2022); Toma & Kikuchi (2024)
Model extraction Example-based Counterfactuals Aïvodji et al. (2020); Kuppa & Le-Khac (2021); Wang et al. (2022)
Feature-based Gradient Milli et al. (2019)
Gradient, Grad-CAM, MASK Yan et al. (2022)
Gradient, Grad-CAM, MASK, LIME Yan et al. (2023b)
Grad-CAM, LIME Yan et al. (2023a)
Gradient, integrated gradients, SmoothGrad (Miura et al., 2024)
@@ 4.3 Unintentional privacy leakage
This subsection discusses unintentional privacy leakage in XAI that occur without malicious intent (Jegorova et al., 2022). Certain leakages can occur due to the natural mechanisms of the training process or through the content of explanations. These may occur during the different AI lifecycle phases or during the course of an explanation (Rawal et al., 2022).
@@ 4.3.1 Training issues
Training issues such as, overfitting and memorization, identified in AI models can lead to privacy leakage. Overfitting is found to aid membership and attribute inference attacks (Yeom et al., 2018). Memorization leads to the model remembering subsets of training data (Song et al., 2017) and occurs during training before overfitting begins (Jegorova et al., 2022). It can cause leakage when data owners deploy models with codebases and training pipelines developed by third parties, such as in MLaaS, allowing sensitive information to be leaked from training data (Song et al., 2017).
@@ 4.3.2 Explanation content
The content of explanations may contain values of sensitive fields. For instance, in example-based explanations such as influence functions, training datapoints potentially containing sensitive fields, are directly revealed to end-users. Karimi et al. (2023) provide another example of unintentional leakage through contrastive explanations, which can lead to inference of sensitive details of individuals whose partial attributes are known. Additionally, interpretable models used as surrogates, can reveal properties of the training data or additional information about the black-box beyond what the model owner intended to share (Blanco-Justicia et al., 2020). In addition to direct content-based leakage, risks may also arise from the inadvertent exposure of explanations to unauthorised users (Kuppa & Le-Khac, 2021). For example, during troubleshooting of error cases, developers or quality engineers may unintentionally access sensitive personal information in the explanation. Moreover, even when direct identifiers are absent, explanations that contain proxy or correlated features can still enable indirect inference.
@@ 5 Privacy Preservation Methods in XAI
To address the privacy risks outlined in Section 4, a growing body of research has emphasized the need for privacy preserving XAI techniques (Aïvodji et al., 2020; Shokri et al., 2021; Zhao et al., 2021). In response to these concerns, several studies have proposed methods to generate explanations while mitigating privacy concerns. Many of these approaches draw upon established principles and methods from the broader domain of privacy preserving ML (PPML), adapting them to specific challenges posed by explanation techniques. This section synthesizes the key contributions in literature to enhance the privacy guarantees of XAI systems in alignment with the objectives of RQ2. We categorize the proposed solutions under the main approaches in PPML. Table LABEL:tab:table3 summarizes these approaches and methods, and Figure 10 presents a consolidated mapping of privacy preserving strategies to specific types of privacy attacks discussed earlier.
@@ 5.1 Differential privacy
Differential privacy (DP) (Dwork et al., 2006) is a widely recognised method that provides a quantifiable definition of privacy and the incremental privacy loss from publishing confidential data (McKay Bowen & Garfinkel, 2021). A mechanism is differentially private if it can hide the participation of any single individual in a dataset (Harder et al., 2020). This can be achieved by using noise and is typically associated with an adverse effect on the accuracy of the system (Harder et al., 2020). By adjusting the privacy budget, $\epsilon$, from 0 to $\infty$, practitioners can manage this trade-off between privacy and accuracy (McKay Bowen & Garfinkel, 2021). Given its robust privacy guarantees, early research in privacy preserving explanations has adopted DP using various strategies to safeguard the training data in interpretable, feature-based, and example-based XAI. In the context of XAI, an explanation is differentially private if it can obscure any single individual in the training dataset (Patel et al., 2022). The technique can be applied at various stages, including the explanation generation algorithm (Patel et al., 2022), the training process of the target model (Cohen & Giryes, 2024; Liu et al., 2024) or directly on the training data (Bozorgpanah & Torra, 2024; Ezzeddine et al., 2024).
@@
Decision trees are popular due to their simplicity and inherent interpretable qualities, however they are prone to privacy leakage (Ferry et al., 2024; Naretto et al., 2022). A number of algorithms have been developed for building decision trees based on DP guarantees (Fletcher & Islam, 2020) offering different trade-offs in privacy and utility. The interpretability of private trees and their subsequent usefulness to XAI users, depend on factors such as the privacy budget per query, tree depth, pruning strategies and termination criteria (Fletcher & Islam, 2020). Nori et al. (2021) applied DP to another type of interpretable model, namely, Explainable Boosting Machines (EBM), to prevent privacy leakage of training data. The resulting privatized system demonstrated good accuracy at low privacy budgets while facilitating correction of errors introduced by noise, the removal of bias and the enforcement of constraints such as monotonicity. Building on this, Baek & Chung (2024) further optimized the utilization of privacy budget in these models to improve accuracy, incorporating techniques such as gradient error optimization and pruning of non-essential features.
@@
Harder et al. (2020) proposed an interpretable model using differentially private locally linear maps with Gaussian mechanism per output class. The filters learned by the model from input images were observed to have higher interpretability compared to feature-based methods, such as integrated gradients and SmoothGrad. However, increasing the number of such maps per class dropped accuracy due to the distribution of privacy budget over additional parameters. In a different approach, Li et al. (2020a) proposed an interpretable model in the form of feedforward-designed convolutional neural network (FF-CNN) made privacy preserving by using DP guarantees on subspace approximation with adjusted bias (Saab). Their findings indicated the the integration of DP was effective in mitigating the risk of reconstruction of input images while maintaining classification accuracy.
@@
For feature-based explanations generated by local linear approximations around the point of interest, Patel et al. (2022) introduced a differentially private approach for loss calculation in the explanation algorithm. The study also proposed an adaptive method of reusing previous explanations for prudent usage of the privacy budget. Nguyen et al. (2023) employed local DP to restrict adversaries from learning the top influential features through aggregated scores in feature-based XAI. While originally proposed to defend against a backdoor security attack exploiting explanations (Severi et al., 2021), the random perturbation of influential features under local DP guarantees was observed to preserve the privacy of those features while maintaining explanation fidelity. Bozorgpanah & Torra (2024) also applied local DP to mask the training dataset and investigated its impact on privacy and utility of feature-based explanations. They introduced an irregularity metric to measure the feature distortion due to privatization of the original dataset and the change in explanation values. Their findings indicate that the use of additive noise on the training dataset caused irregularities, thereby reducing the utility of the explanations. In a related study, Ezzeddine et al. (2024) added calibrated noise to training datasets and evaluated the impact on SHAP explanations using various distance metrics. They observed the change in SHAP values in the privatized systems correlated with the privacy budget and data dependent. Abbasi et al. (2024) used a different approach on the data and employed DP for synthetic data generation for training of different model architectures. They used similarity scores to track the change in explanations while utility loss evaluated the drop in accuracy, thus quantifying the triad of privacy, utility and explainability.
@@
In addition to the aforementioned studies, researchers have also examined the use of DP in the training process of the target model in feature-based (Liu et al., 2024) and example-based (Cohen & Giryes, 2024; Mochaourab et al., 2023) XAI. These investigations have determined mitigation against membership inference attacks for high privacy budgets (Cohen & Giryes, 2024; Liu et al., 2024). The introduction of DP noise serves as a regularization mechanism for target models (Nori et al., 2021) and its mathematical guarantee enables quantification of privacy, making it a compelling choice as a privacy enhancing technology. In the context of XAI, in addition to its application in mitigating membership inference from explanations (Cohen & Giryes, 2024; Liu et al., 2024; Patel et al., 2022), it is also found to mitigate reconstruction of sensitive inputs (Li et al., 2020a). However, the improvement in mitigation of attacks at high privacy budgets and hence degrading accuracy (Cohen & Giryes, 2024; Liu et al., 2024) can be a setback to the use of this technique. In addition to adversely affecting accuracy, its introduction can also deteriorate explanation quality in terms of fidelity (Liu et al., 2024; Patel et al., 2022) with pronounced effects on minority groups (Patel et al., 2022). The technique is also ineffective against attribute inference attacks when there are existing strong correlations between different attributes (Chen et al., 2022a). In such cases, the privacy enhancing benefits of DP may be insufficient to prevent adversaries from inferring sensitive attributes.
@@
Algorithms such as DPSaab (Li et al., 2020a), have been observed to offer a more favourable trade-off between accuracy and privacy. Practitioners can employ strategies such as reusing previously generated differentially private explanations to utilize the privacy budget effectively (Patel et al., 2022). Methods that use local DP, are observed to achieve high faithfulness of explanations with privacy (Nguyen et al., 2023), thus demonstrating that it is possible to balance multiple desirable properties. Hence judicious use of this technique can ensure that privacy is achieved while maintaining reasonable utility of the model and explanations.
@@ 5.2 Cryptography
Cryptographic protocols for privacy preservation in ML use secure algorithms to protect the target model and data. Prominent methods in this category include homomorphic encryption, secret sharing, and secure multi-party computation (Yin et al., 2022). In XAI, these techniques have seen limited application in interpretable and example-based systems. They have also been used in conjunction with other privacy preserving techniques such as federated learning (El Zein et al., 2024; Molhoek & Laanen, 2024; Wu et al., 2020; 2023).
For interpretable tree-based models, Zhao et al. (2023) proposed an additive homomorphic scheme for model owners and query users, for pushing the encrypted model and query data to cloud service providers for inferencing. Adding perturbations to the inference results and query data ensured privacy protection of these assets while maintaining accuracy comparable to non-private inference. In the example-based category, Veugen et al. (2022) proposed a cryptographic method with secure multi-party computation to generate contrastive explanations, while protecting private training data and confidentiality of the model. The algorithm securely trained a binary decision tree to generate fact and foil leaves, which were used as explanations for a query datapoint. Additionally, a synthetic datapoint from the foil leaf was provided to the end-user to enhance explainability.
Cryptographic methods, such as homomorphic encryption, introduce significant computational complexity in the system (Liu et al., 2022a). The use of encryption can deteriorate model transparency, limiting the ability of data scientists to correct errors, inspect data, add features or fine tune the model (Dowlin et al., 2016). Therefore it is essential to implement cryptographic protocols in XAI system components in the right use cases to complement other privacy preservation techniques or when other techniques are infeasible or costly.

### v3

@@ 4.2.4 Property inference
This type of risk pertains to inference of properties from the training data such as global statistics or aggregates (mahloujifarPropertyInferencePoisoning2022), which model owners did not intend on sharing (ganjuPropertyInferenceAttacks2018) (Figure 9). The inferred property may not correspond to features in the training data or be correlated to any feature. A property inference model that uses the model outputs, $f(x)$, to infer such property, $p$, can be expressed as the following:
@@
Membership inference Interpretable Decision tree narettoEvaluatingPrivacyExposure2022
Interpretable (surrogate) Trepan narettoEvaluatingPrivacyExposure2022
Example-based Influence functions shokriExploitingTransparencyMeasures2020; shokriPrivacyRisksModel2021
Counterfactuals kuppaAdversarialXAIMethods2021; pawelczykPrivacyRisksAlgorithmic2023
Self-influence functions cohenMembershipInferenceAttack2024
Feature-based Gradient, integrated gradients, guided backpropagation, LRP, LIME, SmoothGrad shokriExploitingTransparencyMeasures2020; shokriPrivacyRisksModel2021
Integrated gradients, SmoothGrad, VarGrad, Grad-CAM++, SHAP, LIME liuPleaseTellMe2024
Shapley values maLabelOnlyMembershipInference2024
Model inversion Interpretable Decision tree, rule list ferryProbabilisticDatasetReconstruction2024
Example-based Influence functions shokriExploitingTransparencyMeasures2020; shokriPrivacyRisksModel2021
Counterfactuals (native) (goethalsPrivacyIssueCounterfactual2023)
Feature-based Gradient, gradient x input, class activation maps (CAM), Grad-CAM, LRP (zhaoExploitingExplanationsModel2021)
Integrated gradients, DeepLIFT, GradientSHAP, SmoothGrad dudduInferringSensitiveAttributes2022
Shapley values luoFeatureInferenceAttack2022; tomaCombinationsAIModels2024
Model extraction Example-based Counterfactuals aivodjiModelExtractionCounterfactual2020; kuppaAdversarialXAIMethods2021; wangDualCFEfficientModel2022
Feature-based Gradient milliModelReconstructionModel2019
Gradient, Grad-CAM, MASK yanExplainableModelExtraction2022
Gradient, Grad-CAM, MASK, LIME yanExplanationLeaksExplanationguided2023
Grad-CAM, LIME yanExplanationbasedDatafreeModel2023
Gradient, integrated gradients, SmoothGrad (miuraMEGEXDataFreeModel2024)
@@ 4.3 Unintentional privacy leakage
This subsection discusses unintentional privacy leakage in XAI that occur without malicious intent (jegorovaSurveyLeakagePrivacy2022). Certain leakages can occur due to the natural mechanisms of the training process or through the content of explanations. These may occur during the different AI lifecycle phases or during the course of an explanation (rawalRecentAdvancesTrustworthy2022).
@@ 4.3.1 Training issues
Training issues such as, overfitting and memorization, identified in AI models can lead to privacy leakage. Overfitting is found to aid membership and attribute inference attacks (yeomPrivacyRiskMachine2018). Memorization leads to the model remembering subsets of training data (songMachineLearningModels2017) and occurs during training before overfitting begins (jegorovaSurveyLeakagePrivacy2022). It can cause leakage when data owners deploy models with codebases and training pipelines developed by third parties, such as in MLaaS, allowing sensitive information to be leaked from training data (songMachineLearningModels2017).
@@ 4.3.2 Explanation content
The content of explanations may contain values of sensitive fields. For instance, in example-based explanations such as influence functions, training datapoints potentially containing sensitive fields, are directly revealed to end-users. karimiSurveyAlgorithmicRecourse2023 provide another example of unintentional leakage through contrastive explanations, which can lead to inference of sensitive details of individuals whose partial attributes are known. Additionally, interpretable models used as surrogates, can reveal properties of the training data or additional information about the black-box beyond what the model owner intended to share (blanco-justiciaMachineLearningExplainability2020). In addition to direct content-based leakage, risks may also arise from the inadvertent exposure of explanations to unauthorised users (kuppaAdversarialXAIMethods2021). For example, during troubleshooting of error cases, developers or quality engineers may unintentionally access sensitive personal information in the explanation. Moreover, even when direct identifiers are absent, explanations that contain proxy or correlated features can still enable indirect inference.
@@ 5 Privacy Preservation Methods in XAI
To address the privacy risks outlined in Section 4, a growing body of research has emphasized the need for privacy preserving XAI techniques (aivodjiModelExtractionCounterfactual2020; shokriPrivacyRisksModel2021; zhaoExploitingExplanationsModel2021). In response to these concerns, several studies have proposed methods to generate explanations while mitigating privacy concerns. Many of these approaches draw upon established principles and methods from the broader domain of privacy preserving ML (PPML), adapting them to specific challenges posed by explanation techniques. This section synthesizes the key contributions in literature to enhance the privacy guarantees of XAI systems in alignment with the objectives of RQ2. We categorize the proposed solutions under the main approaches in PPML. Table LABEL:tab:table3 summarizes these approaches and methods, and Figure 10 presents a consolidated mapping of privacy preserving strategies to specific types of privacy attacks discussed earlier.
@@ 5.1 Differential privacy
Differential privacy (DP) (dworkOurDataOurselves2006) is a widely recognised method that provides a quantifiable definition of privacy and the incremental privacy loss from publishing confidential data (mckaybowenPhilosophyDifferentialPrivacy2021). A mechanism is differentially private if it can hide the participation of any single individual in a dataset (harderInterpretableDifferentiallyPrivate2020). This can be achieved by using noise and is typically associated with an adverse effect on the accuracy of the system (harderInterpretableDifferentiallyPrivate2020). By adjusting the privacy budget, $\epsilon$, from 0 to $\infty$, practitioners can manage this trade-off between privacy and accuracy (mckaybowenPhilosophyDifferentialPrivacy2021). Given its robust privacy guarantees, early research in privacy preserving explanations has adopted DP using various strategies to safeguard the training data in interpretable, feature-based, and example-based XAI. In the context of XAI, an explanation is differentially private if it can obscure any single individual in the training dataset (patelModelExplanationsDifferential2022). The technique can be applied at various stages, including the explanation generation algorithm (patelModelExplanationsDifferential2022), the training process of the target model (cohenMembershipInferenceAttack2024; liuPleaseTellMe2024) or directly on the training data (bozorgpanahExplainableMachineLearning2024; ezzeddineDifferentialPrivacyAnomaly2024).
@@
Decision trees are popular due to their simplicity and inherent interpretable qualities, however they are prone to privacy leakage (ferryProbabilisticDatasetReconstruction2024; narettoEvaluatingPrivacyExposure2022). A number of algorithms have been developed for building decision trees based on DP guarantees (fletcherDecisionTreeClassification2020) offering different trade-offs in privacy and utility. The interpretability of private trees and their subsequent usefulness to XAI users, depend on factors such as the privacy budget per query, tree depth, pruning strategies and termination criteria (fletcherDecisionTreeClassification2020). noriAccuracyInterpretabilityDifferential2021 applied DP to another type of interpretable model, namely, Explainable Boosting Machines (EBM), to prevent privacy leakage of training data. The resulting privatized system demonstrated good accuracy at low privacy budgets while facilitating correction of errors introduced by noise, the removal of bias and the enforcement of constraints such as monotonicity. Building on this, baekDifferentiallyPrivateExplainable2024 further optimized the utilization of privacy budget in these models to improve accuracy, incorporating techniques such as gradient error optimization and pruning of non-essential features.
@@
harderInterpretableDifferentiallyPrivate2020 proposed an interpretable model using differentially private locally linear maps with Gaussian mechanism per output class. The filters learned by the model from input images were observed to have higher interpretability compared to feature-based methods, such as integrated gradients and SmoothGrad. However, increasing the number of such maps per class dropped accuracy due to the distribution of privacy budget over additional parameters. In a different approach, liDifferentialPrivacyPreservation2020 proposed an interpretable model in the form of feedforward-designed convolutional neural network (FF-CNN) made privacy preserving by using DP guarantees on subspace approximation with adjusted bias (Saab). Their findings indicated the the integration of DP was effective in mitigating the risk of reconstruction of input images while maintaining classification accuracy.
@@
For feature-based explanations generated by local linear approximations around the point of interest, patelModelExplanationsDifferential2022 introduced a differentially private approach for loss calculation in the explanation algorithm. The study also proposed an adaptive method of reusing previous explanations for prudent usage of the privacy budget. nguyenXRandDifferentiallyPrivate2023 employed local DP to restrict adversaries from learning the top influential features through aggregated scores in feature-based XAI. While originally proposed to defend against a backdoor security attack exploiting explanations (severiExplanationGuidedBackdoorPoisoning2021), the random perturbation of influential features under local DP guarantees was observed to preserve the privacy of those features while maintaining explanation fidelity. bozorgpanahExplainableMachineLearning2024 also applied local DP to mask the training dataset and investigated its impact on privacy and utility of feature-based explanations. They introduced an irregularity metric to measure the feature distortion due to privatization of the original dataset and the change in explanation values. Their findings indicate that the use of additive noise on the training dataset caused irregularities, thereby reducing the utility of the explanations. In a related study, ezzeddineDifferentialPrivacyAnomaly2024 added calibrated noise to training datasets and evaluated the impact on SHAP explanations using various distance metrics. They observed the change in SHAP values in the privatized systems correlated with the privacy budget and data dependent. abbasiFurtherInsightsBalancing2024 used a different approach on the data and employed DP for synthetic data generation for training of different model architectures. They used similarity scores to track the change in explanations while utility loss evaluated the drop in accuracy, thus quantifying the triad of privacy, utility and explainability.
@@
In addition to the aforementioned studies, researchers have also examined the use of DP in the training process of the target model in feature-based (liuPleaseTellMe2024) and example-based (cohenMembershipInferenceAttack2024; mochaourabDemonstratorCounterfactualExplanations2023) XAI. These investigations have determined mitigation against membership inference attacks for high privacy budgets (cohenMembershipInferenceAttack2024; liuPleaseTellMe2024). The introduction of DP noise serves as a regularization mechanism for target models (noriAccuracyInterpretabilityDifferential2021) and its mathematical guarantee enables quantification of privacy, making it a compelling choice as a privacy enhancing technology. In the context of XAI, in addition to its application in mitigating membership inference from explanations (cohenMembershipInferenceAttack2024; liuPleaseTellMe2024; patelModelExplanationsDifferential2022), it is also found to mitigate reconstruction of sensitive inputs (liDifferentialPrivacyPreservation2020). However, the improvement in mitigation of attacks at high privacy budgets and hence degrading accuracy (cohenMembershipInferenceAttack2024; liuPleaseTellMe2024) can be a setback to the use of this technique. In addition to adversely affecting accuracy, its introduction can also deteriorate explanation quality in terms of fidelity (liuPleaseTellMe2024; patelModelExplanationsDifferential2022) with pronounced effects on minority groups (patelModelExplanationsDifferential2022). The technique is also ineffective against attribute inference attacks when there are existing strong correlations between different attributes (chenAchievingTransparencyReport2022). In such cases, the privacy enhancing benefits of DP may be insufficient to prevent adversaries from inferring sensitive attributes.
@@
Algorithms such as DPSaab (liDifferentialPrivacyPreservation2020), have been observed to offer a more favourable trade-off between accuracy and privacy. Practitioners can employ strategies such as reusing previously generated differentially private explanations to utilize the privacy budget effectively (patelModelExplanationsDifferential2022). Methods that use local DP, are observed to achieve high faithfulness of explanations with privacy (nguyenXRandDifferentiallyPrivate2023), thus demonstrating that it is possible to balance multiple desirable properties. Hence judicious use of this technique can ensure that privacy is achieved while maintaining reasonable utility of the model and explanations.
@@ 5.2 Cryptography
Cryptographic protocols for privacy preservation in ML use secure algorithms to protect the target model and data. Prominent methods in this category include homomorphic encryption, secret sharing, and secure multi-party computation (yinComprehensiveSurveyPrivacypreserving2022). In XAI, these techniques have seen limited application in interpretable and example-based systems. They have also been used in conjunction with other privacy preserving techniques such as federated learning (elzeinPrivaTreeCollaborativePrivacyPreserving2024; molhoekSecureCounterfactualExplanations2024; wuPrivacyPreservingVertical2020; wuFalconPrivacyPreservingInterpretable2023).
For interpretable tree-based models, zhaoEfficientPrivacypreservingTreebased2023 proposed an additive homomorphic scheme for model owners and query users, for pushing the encrypted model and query data to cloud service providers for inferencing. Adding perturbations to the inference results and query data ensured privacy protection of these assets while maintaining accuracy comparable to non-private inference. In the example-based category, veugenPrivacyPreservingContrastiveExplanations2022 proposed a cryptographic method with secure multi-party computation to generate contrastive explanations, while protecting private training data and confidentiality of the model. The algorithm securely trained a binary decision tree to generate fact and foil leaves, which were used as explanations for a query datapoint. Additionally, a synthetic datapoint from the foil leaf was provided to the end-user to enhance explainability.
Cryptographic methods, such as homomorphic encryption, introduce significant computational complexity in the system (liuWhenMachineLearning2022). The use of encryption can deteriorate model transparency, limiting the ability of data scientists to correct errors, inspect data, add features or fine tune the model (dowlinCryptoNetsApplyingNeural2016). Therefore it is essential to implement cryptographic protocols in XAI system components in the right use cases to complement other privacy preservation techniques or when other techniques are infeasible or costly.